OpenStudy (anonymous):

Let a and b be positive integers such that a^2 = b^3 given 4 divides a, prove that 4 divides b and give an example of positive integers a and b such that a is even and a^2 = b^3, but b is not divisible by 8 please explain each step you did

OpenStudy (anonymous):

a=8, b=4 in this case a^2=b^3 but b is not divisible by 8

OpenStudy (anonymous):

(4)(4)(4) = 64/8 = 8 it is divisible?

OpenStudy (anonymous):

i answered the second part

OpenStudy (anonymous):

yea b is divisible by 8 so this cant be correct

OpenStudy (anonymous):

i cant get it.. 64=b^3 so, b is 4 4 is not divisible by 8

OpenStudy (anonymous):

ahh i see my mistake thanks do you know how to do the first part?

OpenStudy (anonymous):

i am working on that still :/

OpenStudy (anonymous):

alright thanks a lot for your help

OpenStudy (anonymous):

a is divisible by 4 so, a^2 is divisible by 16, that is 2^4 now, a^2=b^3 and b is an integer so, there must be 6 2's when we factorise b^3, otherwise it can't be a cubic number so, there is b^3 is divisible by 2^6=64 so, b is divisible by 4 (proved)

OpenStudy (anonymous):

thanks a lot for your help

OpenStudy (anonymous):

you are most welcome :)